A bullet is fired vertically upward with an initial velocity of 98m/s from the top of a building 100m high. Find,


a) the maximum height reached above the ground


b) the total time before reaching the ground


c) the velocity on landing

1
Expert's answer
2022-02-28T10:40:42-0500

Given:

"v_0=98\:\rm m/s"

"H=100\:\rm m"

a) the maximum height reached above the ground

"H_{\max}=H+\frac{v_0^2}{2g}\\
=100+98^2/(2*9.8)=590\:\rm m"

b) the total time before reaching the ground

"y=-H-v_0t+gt^2\/2"

"0=-100-98t+4.9t^2"

"t=21\:\rm s"

c) the velocity on landing

"v=\sqrt{v_0^2+2gH}\\
=\sqrt{98^2+2*9.8*100}=108\:\rm m/s"